# StreamBridge: Turning Your Offline Video Large Language Model into a Proactive Streaming Assistant

### v1

@@ 1 Introduction
To bridge the gap between offline and streaming video understanding, we introduce StreamBridge, a simple yet effective framework that seamlessly transforms pre-trained offline Video-LLMs into streaming-capable models. In contrast to prior efforts [10; 11; 12; 13], which train streaming models from scratch but fall behind on offline video tasks, StreamBridge leverages the strong generalization capabilities of existing Video-LLMs without requiring full retraining. Concretely, StreamBridge introduces a memory buffer to manage incoming video frames, coupled with a round-decayed compression strategy that merges earlier frame tokens while preserving recent ones, enabling the model to support long-context, multi-turn interactions in streaming scenarios. For proactive capabilities, instead of modifying the base model architecture [14] or introducing streaming-specific objectives [12], both of which can lead to optimization conflicts and issues like probability correction [10], StreamBridge adopts a modular design, by decoupling the proactive capability from the main Video-LLM via a compact activation model. This plug-and-play component operates in parallel with the main Video-LLM, enabling proactive behavior in a flexible and non-intrusive manner while fully preserving the main Video-LLM’s language fluency and general video understanding capabilities.
@@ 2 Related Work
Streaming Video Understanding. Typical tasks in streaming video understanding, such as action recognition [44; 45; 46; 47] and anticipation [48; 49], causally process video inputs using only past and current observations. Recent efforts [50; 12; 13; 51] focus on building Video-LLMs capable of real-time conversation, generating timely responses throughout a live video stream. VideoLLM-Online [10] and Flash-VStream [11] introduce specialized online objectives and memory architectures to handle sequential inputs. MMDuet [14] and ViSpeak [52] add dedicated heads to facilitate proactive response generation. To benchmark streaming video capabilities, several evaluation suites have been proposed, including StreamingBench [21], StreamBench [53], SVBench [54], OmniMMI [55], and OVO-Bench [20]. In contrast to previous approaches that retrain models or tightly couple proactive mechanisms within the backbone, our work leverages the strong generalization abilities of pre-trained offline Video-LLMs [1; 3; 2]. We propose an efficient adaptation framework, combined with a dedicated fine-tuning dataset, to endow these models with streaming capabilities. Furthermore, observing that embedding the activation function into the main model often lead to optimization conflicts and performance degradation [10; 14], we advocate a modular, decoupled design. Our method introduces a compact, plug-and-play activation model that enables proactive behaviors efficiently and non-intrusively.
@@ 3.2.2 Round-Decayed Compression
Online scenarios often involve long, even infinite video streaming, which can lead to significant memory usage and inference latency. Therefore, we propose a round-decayed token compression strategy tailored for multi-turn streaming settings. Specifically, we pre-define a maximum allowable embedding length for the model input. Before each response generation, the model checks whether the current input embedding exceeds . If so, we apply a round-decayed token merging strategy: starting from the earliest dialogue rounds, visual tokens are progressively merged frame-by-frame, until the total length falls below . The merging is implemented via average pooling [56] over adjacent frame tokens. This strategy ensures that the most recent visual context is retained with minimal distortion, thus maintaining the precision of real-time responses while not fully discarding historical visual contexts. At the same time, it significantly improves memory efficiency and reduces inference overhead as in Figure 5. This process is encapsulated in the compression function in Algorithm 1 (line 15). The detailed pseudo codes can be found in Appendix G.
@@ 3.2.3 A Plug-and-play Activation Model
To enable proactive responses, we decouple the activation function from the main Video-LLM. Unlike prior methods that tightly integrate activation mechanisms into the LLM [10; 12; 14; 52], our framework avoids potential interference with the language modeling capacity of the main Video-LLM. Specifically, we propose a parallel pipeline, where a compact external MLLM (e.g., LLaVA-OV-0.5B [3]) is used as an independent activation model, denoted as . As shown in Algorithm 1 (line 9), upon receiving each new frame, the framework simultaneously forwards the current frame (along with the user query and optionally previous frames) to to determine whether a response should be generated. If the activation signal is positive, the buffered embeddings are sent to the LLM for decoding. This design ensures high flexibility and compatibility. Furthermore, in real-time deployment, the , the frame encoder , and the main can run concurrently in parallel threads, enabling more efficient inference.
@@
For training data, we collect a diverse set of temporally annotated video datasets across multiple tasks, including dense video captioning [57; 58], sequential step recognition [59; 60], grounded video question answering [61; 62; 63], temporal video grounding [64], and temporal action detection [65; 66]. For each task, we design specific prompt templates and randomly select among them as during training (details in Appendix A). To supervise activation timing, we insert the response <> at the end of its corresponding annotated timestamp (during inference, the response <> is generated by the larger main Video-LLM). Besides, only the last of frames of each video segment are labeled as positive (i.e., response-worthy), while earlier frames are treated as negatives. is dynamically sampled between 0% and 50% for each training instance, simulating a variety of activation patterns and enhancing the model’s robustness to temporal variations.
@@ 4 Stream-IT Dataset
As analyzed in Section 3.1, streaming scenarios are primarily characterized by multi-turn real-time understanding and proactive responses. However, existing datasets and video sources fall short of fully supporting these requirements [67; 68]. To fill this gap and further enhance the streaming interaction capability of StreamBridge, we introduce Stream-IT—a video-text dataset specifically designed for streaming instruction tuning with an interleaved multi-turn dialogue format.
Datasets for Proactive Understanding. We collect a set of public datasets enriched with timestamp annotations, spanning a wide range of tasks including: (i) Dense Video Captioning [57; 58; 69]; (ii) Sequential Step Recognition [59; 60; 70]; (iii) Grounded VideoQA [61; 71; 72; 73]. All datasets are reformatted into a proactive-style interleaved format: ‘<> <> <>, <> <>, ’, where may be an open-ended query (e.g., “Who is the man going to find?”) or a goal-oriented instruction (e.g., “Show me all the steps for cooking.”). Unlike traditional single-turn datasets where a question is immediately followed by an answer [67; 68], our structure introduces a temporal delay between <> and <> through the inserted video segments <>, simulating proactive response scenarios.
@@ 5.2 Main Results
General Video Understanding. While our StreamBridge is designed for online scenarios, we also verify that it does not downgrade the base model’s performance on standard offline video tasks. As shown in Table 2, models equipped with the StreamBridge framework and fine-tuned on Stream-IT (denoted with ‡) exhibit consistent improvements or maintain comparable performance relative to their original versions. For instance, Oryx-1.5-7B‡ achieves 65.5 on the challenging VideoMME with an increase of 6.7, while LLaVA-OV-7B‡ outperforms its base model across nearly all benchmarks, except LongVideoBench. Likewise, Qwen2-VL-7B‡ achieves competitive results on MVBench, while surpassing its original counterpart on other benchmarks. These outcomes demonstrate that our streaming adaptation enables models to retain, or even exceed their original performance in general video understanding tasks, demonstrating the generality and non-degradability of our method.
@@ B.2 Concatenation Strategy for Constructing StreamingQA-120K
Starting from a pool of 1.28 million filtered short videos sourced from WebVid-10M [19], Panda-70M [18], and InternVid-10M [17], our goal is to iteratively merge semantically similar clips to form long-form video samples. Let denote the entire set of filtered videos. We initiate the process by randomly sampling one clip from as the anchor. We then compute pairwise semantic similarity (based on the middle frame) between and all other videos in . A new clip is sampled according to the similarity distribution (without replacement). The procedure is repeated using as the new anchor, generating from , and so on. This results in a similarity-ordered list of videos . This approach allows for flexible concatenation of any number of clips to construct a long-form sample, by directly selecting a continuous span , without re-computing similarity each time. We also prepare hallucination questions irrelevant to existing video inputs following [20] with a ratio of 0.01%.
@@ Appendix C More Implementation Details
For the activation model, we adopt LLaVA-OV-0.5B as the base model. To improve efficiency, we aggressively pool the frame representations to 16 tokens per frame. During training, only the LoRA adapters, the projector, the score head, and the learnable activation token are trainable. The model is trained for 5 epochs using a fixed learning rate of 2e-5 for the projector, while 2e-4 for the LoRA adapters, score head, and the learnable activation token, with AdamW optimizer. Notably, for both the main VideoLLM and the activation model, we sample frames at 1 FPS to better simulate real-world frame rates. For videos longer than 256 seconds, we uniformly sample 256 frames to fit within the maximum input length constraint. Experiments are conducted on NVIDIA-H100/A100 GPUs. During inference, we sample videos at 2 FPS for short video benchmarks like MVBench, PerpecptionTest, and TempCompass, while 1 FPS for multi-turn real-time understanding benchmarks and long video benchmarks including OVO-Bench, Streaming-Bench, MLVU, LongVideoBench, VideoMME, and EgoSchema.
@@ Appendix D Benchmarks and Metrics
Online Activation Benchmarks. To assess the proactive capabilities of our framework, we evaluate performance on a subset of ET-Bench [64], including Temporal Video Grounding (TVG), Temporal Action Localization (TAL), Dense Video Captioning (DVC), and Sequential Localization Captioning (SLC). These tasks emphasize a shift from passive to active perception, requiring the model to determine when to respond based on upcoming visual inputs, rather than reacting immediately. For example, the Sequential Localization Captioning (SLC) task requires the model to both determine the precise timing of a certain step and output its content. For evaluation metrics, we compute the average F1 score across multiple IoU thresholds (IoU {}) for localization-based tasks. For tasks involving text generation, we adopt sentence-level similarity metrics [83] to measure the semantic alignment between model outputs and ground-truth responses, following prior works [64; 13]. Specifically, the all-MiniLM-L6-v2 model in Sentence-Transformers library is used as the embedding model. Notably, in all these tasks, the question is presented at the beginning of the video, and the model must autonomously decide when to respond. Moreover, the results of TVGF1, TALF1, are the same for our method with different main Video-LLMs, since they use the same activation model and will not be affected by the generated response.
Appendix E Full Results on OVO-Bench and Streaming-Bench
@@
Appendix F Limitations
While our proposed framework and dataset significantly enhance the streaming capabilities of existing offline Video-LLMs, there are still limitations worth noting. First, although the activation model is compact and decoupled from the main LLM, its decision quality is inherently limited by the feature granularity of the small activation model’s visual encoder. In highly complex or ambiguous scenes, the model may still struggle to accurately predict the optimal response timing. Second, while Stream-IT provides large-scale multi-turn, interleaved training data, its construction relies partially on synthetic QA generation and clip concatenation, which, despite careful filtering, may introduce domain shift compared to truly continuous, real-world video streams. Future work could benefit from curating more organically collected long-form streaming videos with naturally evolving events and dialogues. Third, although our round-decayed compression improves inference efficiency and memory usage, it is based on frame-level pruning. Further investigation is needed to develop more adaptive compression strategies that dynamically balance historical preservation and token budget based on scene complexity at a token level. Finally, StreamBridge currently focuses on frame-by-frame streaming under relatively low sampling rates (e.g., 1 FPS). Extending the framework to handle denser frame rates or multi-modal streaming inputs (e.g., audio-visual-text) in real-time remains an important direction for future research.

### v2

@@ 1 Introduction
To bridge the gap between offline and streaming video understanding, we introduce StreamBridge, a simple yet effective framework that seamlessly transforms pre-trained offline Video-LLMs into streaming-capable models. In contrast to prior efforts [10; 11; 12; 13], which train streaming models from scratch but fall behind on offline video tasks, StreamBridge leverages the strong generalization capabilities of existing Video-LLMs without requiring full retraining. This approach allows developers to directly benefit from the rich world knowledge and linguistic fluency of large-scale pre-trained models, while incurring only minimal additional computational cost and data requirements. Concretely, StreamBridge introduces a memory buffer to manage incoming video frames, coupled with a round-decayed compression strategy that merges earlier frame tokens while preserving recent ones, enabling the model to support long-context, multi-modal, and multi-turn interactions in streaming scenarios. For proactive capabilities, instead of modifying the base model architecture [14] or introducing streaming-specific objectives [12], both of which can lead to optimization conflicts and issues like probability correction [10], StreamBridge adopts a modular design, by decoupling the proactive capability from the main Video-LLM via a compact activation model. This plug-and-play component operates in parallel with the main Video-LLM, enabling proactive behavior in a flexible and non-intrusive manner while fully preserving the main Video-LLM’s language fluency and general video understanding capabilities.
@@ 2 Related Work
Streaming Video Understanding. Typical tasks in streaming video understanding, such as action recognition [44; 45; 46; 47] and anticipation [48; 49], causally process video inputs using only past and current observations. Recent efforts [50; 12; 13; 51] focus on building Video-LLMs capable of real-time conversation, generating timely responses throughout a live video stream. VideoLLM-Online [10] and Flash-VStream [11] introduce specialized online objectives and memory architectures to handle sequential inputs. MMDuet [14] and ViSpeak [52] add dedicated heads to facilitate proactive response generation. To benchmark streaming video capabilities, several evaluation suites have been proposed, including StreamingBench [21], StreamBench [53], SVBench [54], OmniMMI [55], and OVO-Bench [20]. In contrast to previous approaches that retrain models or tightly couple proactive mechanisms within the backbone, our work leverages the strong generalization abilities of pre-trained offline Video-LLMs [1; 3; 2]. We propose an efficient adaptation framework, combined with a dedicated fine-tuning dataset, to endow these models with streaming capabilities. Furthermore, observing that embedding the activation function into the main model often lead to optimization conflicts and performance degradation [10; 14], we advocate a modular, decoupled design. Our method introduces a compact, plug-and-play activation model that enables proactive behaviors efficiently and non-intrusively. We also provide additional discussions on how our work relates to the ReKV [56], VideoStreaming [57], and StreamChat [53] in the Appendix F.
@@ 3.2.2 Round-Decayed Compression
Online scenarios often involve long, even infinite video streaming, which can lead to significant memory usage and inference latency. Therefore, we propose a round-decayed token compression strategy tailored for multi-turn streaming settings. Specifically, we pre-define a maximum allowable embedding length for the model input. Before each response generation, the model checks whether the current input embedding exceeds . If so, we apply a round-decayed token merging strategy: starting from the earliest dialogue rounds, visual tokens are progressively merged frame-by-frame, until the total length falls below . The merging is implemented via average pooling [58] over adjacent frame tokens. This strategy ensures that the most recent visual context is retained with minimal distortion, thus maintaining the precision of real-time responses while not fully discarding historical visual contexts. At the same time, it significantly improves memory efficiency and reduces inference overhead as in Figure 5. This process is encapsulated in the compression function in Algorithm 1 (line 15). The detailed pseudo codes can be found in Appendix I.
@@ 3.2.3 A Plug-and-play Activation Model
To enable proactive responses in streaming Video-LLMs, we decouple the activation function from the main Video-LLM. Unlike prior methods that tightly integrate activation mechanisms into the LLM [10; 12; 14; 52], our framework avoids potential interference with the language modeling capacity of the main Video-LLM. Specifically, we propose a parallel pipeline, where a compact external MLLM (e.g., LLaVA-OV-0.5B [3]) is used as an independent activation model, denoted as . As shown in Algorithm 1 (line 9), upon receiving each new frame, the framework simultaneously forwards the current frame (along with the user query and optionally previous frames) to to determine whether a response should be generated. If the activation signal is positive, the buffered embeddings are sent to the LLM for decoding. This design ensures high flexibility and compatibility. Furthermore, in real-time deployment, the , the frame encoder , and the main can run concurrently in parallel threads, enabling more efficient inference.
@@
For training data, we collect a diverse set of temporally annotated video datasets across multiple tasks, including dense video captioning [59; 60], sequential step recognition [61; 62], grounded video question answering [63; 64; 65], temporal video grounding [66], and temporal action detection [67; 68]. For each task, we design specific prompt templates and randomly select among them as during training (details in Appendix A). To supervise activation timing, we insert the response <> at the end of its corresponding annotated timestamp (during inference, the response <> is generated by the larger main Video-LLM). Besides, only the last of frames of each video segment are labeled as positive (i.e., response-worthy), while earlier frames are treated as negatives. is dynamically sampled between 0% and 50% for each training instance, simulating a variety of activation patterns and enhancing the model’s robustness to temporal variations.
@@ 4 Stream-IT Dataset
As analyzed in Section 3.1, streaming scenarios are primarily characterized by multi-turn real-time understanding and proactive responses. However, existing datasets and video sources fall short of fully supporting these requirements [69; 70]. To fill this gap and further enhance the streaming interaction capability of StreamBridge, we introduce Stream-IT—a video-text dataset specifically designed for streaming instruction tuning with an interleaved multi-turn dialogue format.
Datasets for Proactive Understanding. We collect a set of public datasets enriched with timestamp annotations, spanning a wide range of tasks including: (i) Dense Video Captioning [59; 60; 71]; (ii) Sequential Step Recognition [61; 62; 72]; (iii) Grounded VideoQA [63; 73; 74; 75]. All datasets are reformatted into a proactive-style interleaved format: ‘<> <> <>, <> <>, ’, where may be an open-ended query (e.g., “Who is the man going to find?”) or a goal-oriented instruction (e.g., “Show me all the steps for cooking.”). Unlike traditional single-turn datasets where a question is immediately followed by an answer [69; 70], our structure introduces a temporal delay between <> and <> through the inserted video segments <>, simulating proactive response scenarios.
@@ 5.2 Main Results
General Video Understanding. While our method is designed for online scenarios, we also verify that it does not downgrade the base model’s performance on standard offline video tasks. As shown in Table 2, models equipped with the StreamBridge framework and fine-tuned on Stream-IT (denoted with ‡) exhibit consistent improvements or maintain comparable performance relative to their original versions. For instance, Oryx-1.5-7B‡ achieves 65.5 on the challenging VideoMME with an increase of 6.7, while LLaVA-OV-7B‡ outperforms its base model across nearly all benchmarks, except LongVideoBench. Likewise, Qwen2-VL-7B‡ achieves competitive results on MVBench, while surpassing its original counterpart on other benchmarks. These outcomes demonstrate that our streaming adaptation enables models to retain, or even exceed their original performance in general video understanding tasks, demonstrating the generality and non-degradability of our method.
@@ 5.3 In-Depth Analysis
| OVO-Bench | VideoMME | |
| (Real-Time) Avg. | Avg. | |
| 4k | 70.49 | 61.7 |
| 8k | 70.89 | 63.6 |
| 16k | 71.30 | 64.4 |
| 32k | 71.16 | 64.7 |
Impact of . To better understand the impact of , we conducted ablation studies using the Qwen2-VL-StreamBridge model with 1 FPS sampling, varying from 4k to 32k. From Table 6, we observe the following: (1) For streaming tasks (e.g., OVO-Bench Real-Time): Model performance remains relatively stable across varying values, ranging from 70.49% to 71.30%; Accuracy peaks at 16k and slightly declines at 32k, suggesting that further increasing the memory budget yields diminishing returns. This supports our design assumption: in streaming scenarios, models primarily rely on recent context, and older frames can be compressed without significant performance loss. (2) For offline tasks (e.g., VideoMME): Accuracy improves consistently as increases, from 61.7% at 4k to 64.7% at 32k. This means that offline tasks benefit more from retaining the full temporal context and more uncompressed video tokens, especially for long videos that require detailed long-range understanding. StreamBridge can flexibly balance efficiency and performance across both streaming and offline settings by adjusting the memory budget accordingly, and we set = 16k to strike a good balance between them across most tasks.
@@ B.2 Concatenation Strategy for Constructing StreamingQA-120K
Starting from a pool of 1.28 million filtered short videos sourced from WebVid-10M [19], Panda-70M [18], and InternVid-10M [17], our goal is to iteratively merge semantically similar clips to form long-form video samples. Let denote the entire set of filtered videos. We initiate the process by randomly sampling one clip from as the anchor. We then compute pairwise semantic similarity (based on the middle frame) between and all other videos in . A new clip is sampled according to the similarity distribution (without replacement). The procedure is repeated using as the new anchor, generating from , and so on. This results in a similarity-ordered list of videos We formulate this process in Algorithm 2. This approach allows for flexible concatenation of any number of clips to construct a long-form sample, by directly selecting a continuous span , without re-computing similarity each time. We also prepare hallucination questions irrelevant to existing video inputs following [20] with a ratio of 0.01%.
@@ Appendix C More Implementation Details
For the activation model, we adopt LLaVA-OV-0.5B as the base model. To improve efficiency, we aggressively pool the frame representations to 16 tokens per frame. During training, only the LoRA adapters, the projector, the score head, and the learnable activation token are trainable. The model is trained for 5 epochs using a fixed learning rate of 2e-5 for the projector, while 2e-4 for the LoRA adapters, score head, and the learnable activation token, with AdamW optimizer.
Notably, for both the main VideoLLM and the activation model, we sample frames at 1 FPS to better simulate real-world frame rates. For videos longer than 256 seconds, we uniformly sample 256 frames to fit within the maximum input length constraint. Experiments are conducted on NVIDIA-H100/A100 GPUs. During inference, we sample videos at 2 FPS for short video benchmarks like MVBench, PerpecptionTest, and TempCompass, while 1 FPS for multi-turn real-time understanding benchmarks and long video benchmarks including OVO-Bench, Streaming-Bench, MLVU, LongVideoBench, VideoMME, and EgoSchema.
@@ Appendix D Benchmarks and Metrics
Online Activation Benchmarks. To assess the proactive capabilities of our framework, we evaluate performance on a subset of ET-Bench [66], including Temporal Video Grounding (TVG), Temporal Action Localization (TAL), Dense Video Captioning (DVC), and Sequential Localization Captioning (SLC). These tasks emphasize a shift from passive to active perception, requiring the model to determine when to respond based on upcoming visual inputs, rather than reacting immediately. For example, the Sequential Localization Captioning (SLC) task requires the model to both determine the precise timing of a certain step and output its content. For evaluation metrics, we compute the average F1 score across multiple IoU thresholds (IoU {}) for localization-based tasks. For tasks involving text generation, we adopt sentence-level similarity metrics [85] to measure the semantic alignment between model outputs and ground-truth responses, following prior works [66; 13]. Specifically, the all-MiniLM-L6-v2 model in Sentence-Transformers library is used as the embedding model. Notably, in all these tasks, the question is presented at the beginning of the video, and the model must autonomously decide when to respond. Moreover, the results of TVGF1, TALF1, are the same for our method with different main Video-LLMs, since they use the same activation model and will not be affected by the generated response.
Appendix E Broader Impacts
There are many real-world applications of streaming Video-LLMs, such as patient or elderly health monitoring, autonomous driving, and collaborative robots. However, there could be unintended usages and we advocate responsible usage complying with applicable laws and regulations.
Appendix F More Related Works
To address the challenge of long-context understanding in streaming video, several memory and retrieval mechanisms have been proposed. For instance, ReKV [56] introduces a training-free framework that stores and retrieves the Key-Value (KV) caches of processed frames, enabling offline models to answer user queries efficiently by reloading only the most relevant context. Besides, VideoStreaming [57] employs a memory-propagated encoding architecture where a condensed representation of the preceding clip serves as historical context for encoding the next, combined with an adaptive selection of memories for question-answering. Moreover, StreamChat [53] proposes a hierarchical memory system comprising short-term, long-term, and dialogue components to facilitate complex streaming interactions, and also contributes the StreamBench benchmark for evaluating diverse streaming scenarios. While these methods effectively advance long-context retention for reactive question-answering, StreamBridge differs by introducing a round-decayed compression strategy specifically tailored for multi-turn real-time interactions, which efficiently prunes redundant historical tokens while preserving recent context with high fidelity. Moreover, StreamBridge introduces a decoupled, lightweight activation model. This plug-and-play component operates in parallel with the main Video-LLM, enabling continuous proactive responses. These designs, supported by our dedicated Stream-IT dataset, effectively transform general-purpose offline models into versatile and proactive streaming assistants without compromising their core performance.
Appendix G Limitations
Although our proposed framework and dataset significantly enhance the streaming capabilities of existing offline Video-LLMs, there are still limitations worth noting. First, while Stream-IT provides large-scale multi-turn, interleaved training data, its construction relies partially on synthetic QA generation and clip concatenation, which, despite careful filtering, may introduce domain shift compared to truly continuous, real-world video streams. Future work could benefit from curating more organically collected long-form streaming videos with naturally evolving events and dialogues. Second, StreamBridge currently focuses on frame-by-frame streaming under relatively low sampling rates (e.g., 1 FPS). Extending the framework to handle denser frame rates or multi-modal streaming inputs (e.g., audio-visual-text) in real-time remains an important direction for future research.
Appendix H Full Results on OVO-Bench and Streaming-Bench